OpenStudy (anonymous):

Find the derivative of f(x) = -2x2 + 11x at x = 9

OpenStudy (anonymous):

-4x+11=-36+11=25

OpenStudy (anonymous):

-4x+11

OpenStudy (ddcamp):

f'(x) = -4x + 11 f'(9) = -36 + 11 = -25

OpenStudy (anonymous):

These are the choices -25 -243 63 19

OpenStudy (anonymous):

at x=9 it's -36+11 =-25

OpenStudy (anonymous):

-25

OpenStudy (anonymous):

-25